fred blended coffee worth 95.00 pesos per kg with coffee worth 115.00 pesos per kg to make 50kg of coffee that will be sold at 107.00 pesos per kg . how many kilograms of each kind did he blend?
Answer by ewatrrr(24785)

BLEND
115x + 95(50kg-x) = 107(50kg)
20x = 12(50kg)
x = 12(50kg)/20 = 30kg, amount of the 115pesos coffee